OpenStudy (lilsis76):

find all the real solutions. Are the real solutions the answers? 12x^2 + 2x -2 = 0

OpenStudy (anonymous):

please wait...i will give you the answers...

OpenStudy (lilsis76):

okay, thank you. I just dont know where or how to start the steps to solving. I have a very hard time with doing the first term higher than 2x^2. that is where i start panicking.

OpenStudy (debbieg):

You have a common factor of 2 that you can divide through by. Then it will factor. Once you have it factored, you just get one solution from each of the linear factors. :) Don't panic, this is not that hard!

OpenStudy (lilsis76):

haha I know its not hard, but the way I see it looks complicated. okay then I have this: \[2(6x ^{2}+x-1)\] @DebbieG

OpenStudy (debbieg):

\[\Large 12x^2 + 2x -2 =0\]\[\Large 6x^2 + x -2 =0\]

OpenStudy (debbieg):

Excellent! Now... it's an equation, so you can divide both sides by that 2, and you have what I show above. That's an EQUIVALENT EQUATION to what you started with, so has the same solutions, so now you just have to solve THAT. Do you know how to solve it? HINT: If that left side FACTORS, that's usually the easiest way to go!

OpenStudy (debbieg):

oops, wait, my last term should be -1, lol

OpenStudy (debbieg):

you had it right though :)\[\Large 6x^2 + x -1 =0\]

OpenStudy (anonymous):

do you have the answer now?

OpenStudy (anonymous):

I found the answer!

OpenStudy (lilsis76):

SORRY GUYS! I lost connection. Im back

OpenStudy (debbieg):

welcome back :)

OpenStudy (lilsis76):

thank you Debbie :) okay so i had that answer you had 2(6x^2 +x -1) =0

OpenStudy (lilsis76):

how can i factor the inside if there is a 6x^2?? I got...hold on let me see what I got

OpenStudy (debbieg):

Right, and then just divide through by that pesky 2, we don't need it. So the equation becomes:\[\Large 6x^2 + x -1 =0\]

OpenStudy (lilsis76):

i got a -2 and +3 from factoring, is this correct?

OpenStudy (debbieg):

Do some trial and error. You need a product of 6(-1)=-6 that sumes to +1.

OpenStudy (debbieg):

*sum

OpenStudy (anonymous):

\[\Delta \] @DebbieG you want to continue with Delta?

OpenStudy (debbieg):

Show your factored form of the trinomial, what did you get? (.........)(........)

OpenStudy (lilsis76):

how did you get a +1? I dont think Delta, I dont know what that is

OpenStudy (debbieg):

I don't know what you mean either, @PFEH.1999 . Just need to factor that left hand side.

OpenStudy (debbieg):

Crud, sorry, I meant sums to -1.... LOL!

OpenStudy (lilsis76):